Circular No. 8824

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2615 OPHIUCHI = NOVA OPHIUCHI 2007 S. Nakano, Sumoto, Japan, reports the discovery by Hideo Nishimura

(Miyawaki, Kakegawa, Shizuoka-ken) of a possible nova (mag 10.2) on two Kodak T-Max 400 films (limiting mag 12) taken on Mar. 19.812 UT using a Pentax 6×7 camera (+ 200-mm f/4 lens), the new object’s position given as α = 17^{h}42^{m}49^{s}, δ = −23^{o}40′07′′ (equinox 2000.0). Nothing was visible at this location on a film taken by Nishimura on Mar. 17.82 and earlier survey films since 2005 (limiting mag 11.5–12). Nakano adds that K. Kadota (Ageo, Japan) confirms the new star at mag 9.8 on an unfiltered CCD exposure taken on Mar. 20.756 with a 0.25-m f/5 reflector, at position end figures 44^{s}.00, 35′′.1; Kadota adds that there is no USNO-A2.0 star within 0′.1 of this position (the closest star having position end figures 43^{s}.41, 34′′.9 and blue mag 18.5, red mag 16.8). Also, K. Itagaki has confirmed the new star using 0.30-m f/7.8 reflector at his Takanezawa station, his CCD image on Mar. 20.806 yielding mag 9.7 and position end figures 43^{s}.99, 35′′.0. H. Yamaoka, Kyushu University, reports the apparently independent discovery of this apparent nova by Yuji Nakamura (Kameyama, Mie, Japan) at mag 10.0 on unfiltered digital-camera images taken on Mar. 20.812 UT. The variable’s position end figures were given by Nakamura as 44^{s}.5, 16′′ (uncertainty ± 20′′); nothing was detected at this position on an image taken by Nakamura on Mar. 17.825 (limiting mag ∼ 13.0). Yamaoka notes that a star of mag 12-13 is located 1′ northeast of the new object. Nakano further writes that, following announcement of this nova on CBET 900, he received word of an apparently independent discovery at mag 10.2 by A. Tago (Tsuyama, Okayama-ken, Canon 20Da digital camera + 70-mm f/3.2 lens) on three frames taken around Mar. 20.8; Tago’s images from Mar. 18.82 show nothing at the nova’s position to mag 11.3.

Following posting on the Central Bureau’s unconfirmed-objects webpage, H. Naito and S. Narusawa, Nishi-Harima Astronomical Observatory, report that a low-resolution spectrogram (range 410–670 nm; resolution 1500 at Hα) of this possible nova, obtained on Mar. 20.84 UT with the 2.0-m NAYUTA telescope (+ MALLS), shows Balmer-lines having P-Cyg features and Fe II lines (multiplets 42, 49, 74), suggesting that the variable is a “Fe II”-type nova. The FWHMs of the Hα, Hβ, and Hγ emissions are 920, 810, and 790 km/s, and the displacement of the P-Cyg absorptions from the Balmer emission peaks is 940, 820, and 830 km/s, respectively.

N. N. Samus, Institute of Astronomy, Russian Academy of Sciences, informs us that this nova has been given the designation V2615 Oph.

2007 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green